Session: Feb 2000 (session 1)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (mauro at Thu Mar 2 00:10:04 MET 2000) Success
All scan recorded. Sky is cloudy and fog GPS-Formatter = - 67.1 uSec

Jodrell (Alastair Gunn at Thu Mar 2 05:39:16 MET 2000) Minor failures
Telescope control lost between 14:10 and 14:30 UT. Telescope parked due to high winds at about 15:46 UT but back on again at about 16:17 UT. No other known problems.

Cambridge (Alastair Gunn at Thu Mar 2 05:39:43 MET 2000) Minor failures
Frequent FS crashes caused by MERLIN control (at 06:36, 07:37, 14:10, 15:38, 16:31, 17:38, 17:51 and 18:30). Some scans missed or short due to these problems.

Shanghai (Huang Xinyong at Thu Mar 2 06:43:42 MET 2000) Success
The second tape was used for fourth tape again. So, the data of some scans on second tape (pass=8, footage from 12800 to 4000) are covered.

Shanghai (Huang Xinyong at Thu Mar 2 06:50:09 MET 2000) Success
The second tape was used for fourth tape again. So, the data of some scans on second tape (pass=8, 12800 to 4000) are covered.

Torun (K.Borkowski at Thu Mar 2 09:28:01 MET 2000) Success
Observations off source between 11:28 and 11:38 because of cable unwrap slewing in the azimuth. Otherwise no known problems.

Effelsberg (A. Kraus at Thu Mar 2 10:55:25 MET 2000) Minor failures
Stopped from UT1601-1643 because of snowfall. Several scans had been skipped for flux density measurements (as requested by the PI).

Noto (Noto VLBI friend at Thu Mar 2 12:21:00 MET 2000) Minor failures
Problems with azimuth tracking. Lost observations from 12:54 to 13:13 UT

Onsala (A. Polatidis at Thu Mar 2 12:56:43 MET 2000) Minor failures
Observations started at 0823, PASS 04R. The antenna was stowed during the night due to a hardware problem which also affected part of the previous experiment. Observations resumed after the engineers fixed the problem. There were strong winds during most of the night so we would not be observing anyway. We will send 3 tapes to the correlator.

Westerbork (Tony Foley at Thu Mar 2 15:55:56 MET 2000) Minor failures
Scans from 23:45 to 03:20 out of hour angle (HA) range, so were replaced by local calibration. Vacuum loss on the tape caused loss of scans from 07:15 to 08:20. However we had to replace scans from 07:35 to 08:21 anyway as they were also outside HA range. Phase stability was poor, often increasing SEFD of the array by ~40%.

Calibration (Alastair Gunn at Wed Apr 19 13:10:59 METDST 2000) Success
Calibration data available from vlbeer: see file feb00/er011.antab for details